To determine the eigenvalues and eigenvectors of the following matrix :

2 -2 0

P = -2 1 -2

0 -2 0


1
Expert's answer
2021-02-24T07:29:38-0500

"P=\left(\begin{array}{rrr}2&-2&0\\-2&1&-2\\0&-2&0\end{array}\right)"


Eigenvalues "\lambda_1", "\lambda_2", "\lambda_3" are the roots of the characteristic polynomial of "P" .

Obtain the characteristic polynomial taking the determinant


"\left|\begin{array}{rrr}2-\lambda&-2&0\\-2&1-\lambda&-2\\0&-2&-\lambda\end{array}\right|="


"(2-\lambda)(1-\lambda)(-\lambda)+(-2)\cdot(-2)\cdot0+" "0\cdot(-2)\cdot(-2)-"

"0\cdot(1-\lambda)\cdot0-(-2)\cdot(-2)\cdot(-\lambda)-" "(2-\lambda)\cdot(-2)\cdot(-2)="


"(2-\lambda-2\lambda+\lambda^2)(-\lambda)+0+0-0+4\lambda-(2-\lambda)\cdot4="


"(2-3\lambda+\lambda^2)(-\lambda)+4\lambda-8+4\lambda="


"-2\lambda+3\lambda^2-\lambda^3+8\lambda-8="


"-\lambda^3+3\lambda^2+6\lambda-8"


Find the roots factoring the polynomial


"-(\lambda^3-3\lambda^2-6\lambda+8)="


"-(\lambda^3+8-3\lambda^2-6\lambda)="


"-((\lambda^3+8)-(3\lambda^2+6\lambda))="


"-((\lambda^3+2^3)-3\lambda(\lambda+2))="


"-((\lambda+2)(\lambda^2-2\lambda+2^2)-3\lambda(\lambda+2))="


"-((\lambda+2)(\lambda^2-2\lambda+4)-3\lambda(\lambda+2))="


"-(\lambda+2)(\lambda^2-2\lambda+4-3\lambda)="


"-(\lambda+2)(\lambda^2-5\lambda+4)="


"-(\lambda+2)(\lambda^2-4\lambda-\lambda+4)="


"-(\lambda+2)((\lambda^2-4\lambda)-(\lambda-4))="


"-(\lambda+2)(\lambda(\lambda-4)-(\lambda-4))="


"-(\lambda+2)(\lambda-1)(\lambda-4)"


So the roots of polynomial are


"\lambda_1=-2",

"\lambda_2=1",

"\lambda_3=4".


Find eigenvectors solving the homogeneous system substituting "\lambda" by "-2", "1", "4"


"\left\{\begin{array}{rrrr}(2-\lambda)x_1&-2x_2&&=0\\-2x_1&+(1-\lambda)x_2&-2x_3&=0\\&-2x_2&-\lambda x_3&=0\end{array}\right."


"\lambda=\lambda_1=-2"


"\left\{\begin{array}{rrrr}4x_1&-2x_2&&=0\\-2x_1&+3x_2&-2x_3&=0\\&-2x_2&+2x_3&=0\end{array}\right."

Use Gaussian elimination modifying the augmented matrix

"\left(\begin{array}{rrr}4&-2&0\\-2&3&-2\\0&-2&2\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)\sim" "L_2+\frac12L_1\to L_2\sim"


"\left(\begin{array}{lll}4&-2&0\\-2+\frac12\cdot4&3+\frac12\cdot(-2)&-2+\frac12\cdot0\\0&-2&2\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)="


"\left(\begin{array}{rrr}4&-2&0\\0&2&-2\\0&-2&2\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)\sim" "L_3+L_2\to L_3\sim"


"\left(\begin{array}{rrr}4&-2&0\\0&2&-2\\0+0&-2+2&2-2\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)="


"\left(\begin{array}{rrr}4&-2&0\\0&2&-2\\0&0&0\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)".

System represented as the matryx is

"\left\{\begin{array}{rrrr}4x_1&-2x_2&&=0\\&+2x_2&-2x_3&=0\end{array}\right."

Let "x_3" be free while "x_1" and "x_2" be dependent on "x_3" . Then

"\left\{\begin{array}{rrrr}4x_1&&&=2x_2\\&2x_2&&=2x_3\end{array}\right.\Rightarrow" "\left\{\begin{array}{rrrr}4x_1&&&=2x_3\\&x_2&&=x_3\end{array}\right.\Rightarrow" "\left\{\begin{array}{rrrr}x_1&&&=\frac12x_3\\&x_2&&=x_3\end{array}\right."

Thus, the eigenvector corresponding to the eigenvalue "\lambda_1=-2" is

"\bold{v}_1=\left(\begin{array}{r}x_1\\x_2\\x_3\end{array}\right)=" "\left(\begin{array}{r}\frac12x_3\\x_3\\x_3\end{array}\right)".

Here "x_3" is allowed to take any value. In particular, if "x_3=1"

"\bold{v}_1=\left(\begin{array}{r}\frac12\\1\\1\end{array}\right)".


"\lambda=\lambda_2=1"


"\left\{\begin{array}{rrrr}x_1&-2x_2&&=0\\-2x_1&&-2x_3&=0\\&-2x_2&-x_3&=0\end{array}\right."

Use Gaussian elimination modifying the augmented matrix

"\left(\begin{array}{rrr}1&-2&0\\-2&0&-2\\0&-2&-1\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)\sim" "L_2+2L_1\to L_2\sim"


"\left(\begin{array}{lll}1&-2&0\\-2+2\cdot1&0+2\cdot(-2)&-2+
2\cdot0\\0&-2&-1\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)="


"\left(\begin{array}{rrr}1&-2&0\\0&-4&-2\\0&-2&-1\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)\sim" "L_3-\frac12L_2\to L_3\sim"


"\left(\begin{array}{lll}1&-2&0\\0&-4&-2\\0-\frac12\cdot0&-2-\frac12\cdot(-4)&-1-\frac12\cdot(-2)\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)="


"\left(\begin{array}{rrr}1&-2&0\\0&-4&-2\\0&0&0\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)"

System represented as the matryx is

"\left\{\begin{array}{rrrr}x_1&-2x_2&&=0\\&-4x_2&-2x_3&=0\end{array}\right."

Let "x_3" be free while "x_1" and "x_2" be dependent on "x_3" . Then

"\left\{\begin{array}{rrrr}x_1&&&=2x_2\\&-4x_2&&=2x_3\end{array}\right.\Rightarrow" "\left\{\begin{array}{rrrl}x_1&&&=2\left(-\frac12x_3\right)\\&x_2&&=-\frac12x_3\end{array}\right.\Rightarrow" "\left\{\begin{array}{rrrl}x_1&&&=-x_3\\&x_2&&=-\frac12x_3\end{array}\right."

Thus, the eigenvector corresponding to the eigenvalue "\lambda_2=1" is

"\bold{v}_2=\left(\begin{array}{r}x_1\\x_2\\x_3\end{array}\right)=" "\left(\begin{array}{r}-x_3\\-\frac12x_3\\x_3\end{array}\right)".

Here "x_3" is allowed to take any value. In particular, if "x_3=1"

"\bold{v}_2=\left(\begin{array}{r}-1\\-\frac12\\1\end{array}\right)".


"\lambda=\lambda_3=4"


"\left\{\begin{array}{rrrr}-2x_1&-2x_2&&=0\\-2x_1&-3x_2&-2x_3&=0\\&-2x_2&-4x_3&=0\end{array}\right."

Use Gaussian elimination modifying the augmented matrix

"\left(\begin{array}{rrr}-2&-2&0\\-2&-3&-2\\0&-2&-4\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)\sim" "L_2-L_1\to L_2\sim"


"\left(\begin{array}{lll}-2&-2&0\\-2-(-2)&-3-(-2)&-2-0\\0&-2&-4\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)="


"\left(\begin{array}{rrr}-2&-2&0\\0&-1&-2\\0&-2&-4\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)\sim" "L_3-2L_2\to L_3\sim"


"\left(\begin{array}{lll}-2&-2&0\\0&-1&-2\\0-2\cdot0&-2-2\cdot(-1)&-4-2\cdot(-2)\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)="


"\left(\begin{array}{rrr}-2&-2&0\\0&-1&-2\\0&0&0\end{array}\right|\left.\begin{array}{r}0\\0\\0\end{array}\right)" .

System represented as the matryx is

"\left\{\begin{array}{rrrr}-2x_1&-2x_2&&=0\\&-x_2&-2x_3&=0\end{array}\right."

Let "x_3" be free while "x_1" and "x_2" be dependent on "x_3" . Then

"\left\{\begin{array}{rrrr}-2x_1&&&=2x_2\\&-x_2&&=2x_3\end{array}\right.\Rightarrow" "\left\{\begin{array}{rrrl}x_1&&&=-x_2\\&x_2&&=-2x_3\end{array}\right.\Rightarrow" "\left\{\begin{array}{rrrl}x_1&&&=2x_3\\&x_2&&=-2x_3\end{array}\right."

Thus, the eigenvector corresponding to the eigenvalue "\lambda_2=1" is

"\bold{v}_3=\left(\begin{array}{r}x_1\\x_2\\x_3\end{array}\right)=" "\left(\begin{array}{r}2x_3\\-2x_3\\x_3\end{array}\right)".

Here "x_3" is allowed to take any value. In particular, if "x_3=1"

"\bold{v}_3=\left(\begin{array}{r}2\\-2\\1\end{array}\right)".



Answer:

Eigenvalues

"\lambda_1=-2", "\lambda_2=1", "\lambda_3=4".

Eigenvectors

"\bold{v}_1=\left(\begin{array}{r}\frac12\\1\\1\end{array}\right)t_1", "\bold{v}_2=\left(\begin{array}{r}-1\\-\frac12\\1\end{array}\right)t_2", "\bold{v}_3=\left(\begin{array}{r}2\\-2\\1\end{array}\right)t_3",

where factors "t_1" , "t_2" , "t_3" may take any value.